Assume that the following data has been determined for the development and sale of a new digital thermometer for home use.
Development cost $480,000,
Sales volume 20,000 units per year,
Sales lifetime 3 years,
Production costs per unit $5

The sale price necessary to break even within two years is:

A) $6

B) $11

C) $5

D) $17

E) none of the other answers are correct
Answer #1

SALES VOLUME PER YEAR = 20000

BREAK EVEN NEEDED IN 2 YEARS, SO TOTAL BREAK EVEN SALES = 20000 X 2 = 40000 UNITS

BREAK EVEN POINT = DEVELOPMENT COST/(SELLING PRICE - VARIABLE COST)

40000 = 480000/(SELLING PRICE - 5)

SELLING PRICE - 5 = 480000/40000 = 12

SELLING PRICE = 17

ANSWER : D: $17 (Thumbs up please)